RETURN ON EQUITY

Pacific Packaging's ROE last year was only 2%; but its management has developed a new operating plan that calls for a debt-to-capital ratio of 55%, which will result in annual interest charges of $760,000. The firm has no plans to use preferred stock and total assets equal total invested capital. Management projects an EBIT of $1,862,000 on sales of $19,000,000, and it expects to have a total assets turnover ratio of 2.2. Under these conditions, the tax rate will be 40%. If the changes are made, what will be the company's return on equity? Do not round intermediate calculations. Round your answer to two decimal places.

%
Answer #1

Solution:-

Asset turnover ratio= Sales/Total assets

Total assets= Sales/total assets turnover ratio= 19,000,000/2.2= $8,636,363.6m

We know total assets is equal to invested capital. Therefore,

Equity shareholders funds= Total assets * (1-debt to capital ratio)= $8,636,363.6*(1-0.55)= $3,886,363.6m

Now, earnings for equity investors= (EBIT - interest)*(1-tax rate)= (18,620,00-760,000)*(1-0.4)= $661,200

Therefore, ROE= Earnings for equity investors/Equity shareholders funds= $661,200/$3,886,363.6= 17.01%